# Uncertainty Estimation and Quantification for LLMs: A Simple Supervised Approach

### v1

@@ 3.4 Three regimes of supervised uncertainty estimation
A natural implementation of our supervised learning approach involves using an LLM to generate the response $\bm{y}$ for input $\bm{x}$, and extracting insights on confidence from the same LLM’s hidden layers’ activations. This method functions effectively with white-box LLMs where hidden activations are accessible, but with black-box LLMs, which restrict access to hidden activations, alternative black-box uncertainty estimation methods become necessary.
We observe that obtaining hidden layers’ activations merely requires an LLM and the prompt-response pair $(\bm{x},\bm{y})$. Therefore, it is not mandatory for $(\bm{x},\bm{y})$ to be generated by the LLM that provides the hidden layers’ activations. Based on this observation, we note that the extra knowledge of uncertainty can come from the hidden layers of any white-box LLM that takes as input the $(\bm{x},\bm{y})$ pair, not necessarily from the LLM that generates $(\bm{x},\bm{y})$. It is natural to start with the belief that the LLM which generates $(\bm{x},\bm{y})$ should have more information about the uncertainty of $\bm{y}$. However, any white-box LLM can output the hidden activations corresponding to $(\bm{x},\bm{y})$.
To clarify further, once $(\bm{x},\bm{y})$ is generated by the LLM $p_{\theta}$, all of $p_{\theta}$’s knowledge about the uncertainty of $\bm{y}$ is encoded within its internal representations. However, another LLM $q_{\theta}$ can also evaluate $(\bm{x},\bm{y})$, with its understanding of uncertainty reflected in $q_{\theta}$’s internal representations as well.
@@
Grey-box supervised uncertainty estimation (Gb-S): This Gb-S regime constructs the features only from the grey-box source, that is, those features relying on the probability and the entropy (such as those in Table 5 in Appendix A.1) but it ignores the hidden-layer activations.
@@ 4.1 LLMs, tasks, benchmarks, and performance metrics
LLMs. For our numerical experiments, we mainly consider two open-source LLMs, LLaMA2-7B (Touvron et al., 2023) and Gemma-7B (Gemma Team et al., 2024) as $p_{\theta}$ defined in Section 2. For certain experiments, we also employ the models of LLaMA2-13B and Gemma-2B. We also use their respective tokenizers as provided by Hugging Face. We do not change the parameters/weights $\theta$ of these LLMs.
Tasks and Datasets. We mainly consider three tasks for uncertainty estimation, question answering, multiple choice, and machine translation. All the labeled datasets for these tasks are in the form of $\{(\bm{x}_{i},\bm{y}_{i,\text{true}})\}_{i=1}^{n}$ where $\bm{x}_{i}$ can be viewed as the prompt for the $i$-th sample and $\bm{y}_{i,\text{true}}$ the true response. We adopt the few-shot prompting when generating the LLM’s response $\bm{y}_{i}$, and we use 5 examples in the prompt of the multiple-choice task and 3 examples for the remaining natural language generation tasks. This enables the LLM’s in-context learning ability (Radford et al., 2019; Zhang et al., 2023) and ensures the LLM’s responses are in a desirable format. We defer more details of the few-shot prompting to Appendix A.2. The three tasks are:
@@
  • Multiple choice. We consider the Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al., 2020), a collection of 15,858 questions covering 57 subjects across STEM. Due to the special structure of the dataset, the generated output $\bm{y}_{i}$ and the correct answer $\bm{y}_{\text{true},i}\in\{\text{A, B, C, D}\}$. Therefore, this task can also be regarded as a classification problem for the LLM by answering the question with one of the four candidate choices.
@@
Benchmarks. We compare our approach with a number of the state-of-the-art benchmarks for the problem. Manakul et al. (2023) give a comprehensive survey of the existing methods and compare four distinct measures for predicting sentence generation uncertainty. The measures are based on either the maximum or average values of entropy or probability across the sentence, including Max Likelihood, Avg Likelihood, Max Ent, and Avg Ent defined in Table 5. We note that each of these measures can be applied as a single uncertainty estimator, and they are all applied in an unsupervised manner that does not require additional supervised training. In particular, in applying these measures for the MMLU dataset, since the answer only contains one token from $\{\text{A, B, C, D}\}$, we use the probabilities and the entropy (over these four tokens) as the benchmarks which represent the probability of the most likely choice and the entropy of all choices, respectively. Kuhn et al. (2023) generate multiple answers, compute their entropy in a semantic sense, and define the quantity as semantic entropy. This semantic-entropy uncertainty (SU) thus can be used as an uncertainty estimator for the LLM’s responses. Tian et al. (2023) propose the approach of asking the LLM for its confidence (denoted as A4U) which directly obtains the uncertainty score from the LLM itself.
Our methods. We follow the discussions in Section 3.4 and implement three versions of our proposed supervised approach: black-box supervised (Bb-S), grey-box supervised (Gb-S), and white-box supervised (Wb-S). These models have the same pipeline of training the uncertainty estimation model and the difference is only on the availability of the LLM. For the Bb-S method, we use the Gemma-7B as the model $q_{\theta}$ to evaluate the uncertainty of LLaMA-7B $p_{\theta}$ (treated as a black-box), and reversely, use LLaMA-7B to evaluate Gemma-7B. The supervised uncertainty model $\hat{g}$ is trained based on the random forest model (Breiman, 2001). Details on the feature construction and the training of the random forest model are deferred to Appendix A.3.
Performance metrics. For the model evaluation, we follow Filos et al. (2019); Kuhn et al. (2023) and compare the performance of our methods against the benchmark using the generated uncertainty score to predict whether the answer is correct. The area under the receiver operator characteristic curve (AUROC) metric is employed to measure the performance of the uncertainty estimation. As discussed in Section 3.2, AUROC works as a good metric for the uncertainty estimation task whereas for the uncertainty calibration task, we follow the more standard calibration metrics and present the results in Section 4.4.
@@ 4.2.1 Question answering and machine translation
Max Pro Avg Pro Max Ent Avg Ent SU A4C Bb-S Gb-S Wb-S
TriviaQA G-7B 0.797 0.769 0.783 0.759 0.712 0.512 0.964 0.893 0.963
L-7B 0.835 0.826 0.833 0.808 0.817 0.528 0.856 0.852 0.879
CoQA G-7B 0.720 0.701 0.716 0.671 0.657 0.509 0.747 0.729 0.769
L-7B 0.699 0.672 0.692 0.644 0.647 0.508 0.706 0.713 0.751
WMT-14 G-7B 0.735 0.813 0.507 0.814 0.536 0.590 0.730 0.822 0.828
L-7B 0.623 0.700 0.588 0.689 0.509 0.517 0.654 0.708 0.752
@@
Table 1: Out-of-sample AUROC performance for benchmarks and our methods on natural language generation tasks. G-7B and L-7B represent Gemma-7B and LLaMA2-7B, respectively. The columns Max Pro, Avg Pro, Max Ent, and Avg Ent all come from Manakul et al. (2023). The column SU implements the semantic uncertainty estimation by Kuhn et al. (2023), and the column A4C implements the ask-for-confidence method by Tian et al. (2023). The columns Bb-S, Gb-S, and Wb-S represent respectively the three regimes (black-box supervised, grey-box supervised, and white-box supervised) of our supervised method with details in Section 3.4.
@@
We make several remarks on the numerical results. First, our methods generally have a better performance than the existing benchmarks. Note that the existing benchmarks are mainly unsupervised and based on one single score, and also that our method proceeds with the most standard pipeline for supervised training of an uncertainty estimation model. The advantage of our method should be attributed to the supervised nature and the labeled dataset. While these unsupervised benchmark methods can work in a larger scope than these NLP tasks (though they have not been extensively tested on open questions yet), our methods rely on the labeled dataset. But in addition to these better numbers, the experiment results show the potential of labeled datasets for understanding the uncertainty in LLM’s responses. In particular, our method Gb-S uses the exactly same features as the benchmark methods, and it shows that some minor supervised training can improve a lot upon the ad-hoc uncertainty estimation based on one single score such as Max Pro or Max Ent.
@@
Second, our method Wb-S has a clear advantage over our method Gb-S. Note that these two methods differ in that the Wb-S uses the hidden activations while the Gb-S only uses probability-related (and entropy-related) features. This implies that the hidden activations do contain uncertainty information which we will investigate more in the next subsection. Also, we note from the table that there is no single unsupervised grey-box method (under the Benchmarks columns) that consistently surpasses others across different datasets/NLP tasks. For example, among all these unsupervised benchmark methods, Avg Ent emerges as a top-performing one for the Gemma-7B model when applied to the machine translation task, but it shows the poorest performance for the same Gemma-7B model when tested on the question-answering TriviaQA dataset. This inconsistency highlights some caveats when using the unsupervised approach for uncertainty estimation of LLMs.
@@
Lastly, we note that the Bb-S method has a similar or even better performance as the Wb-S method. As discussed in Section 3.4, the performance of uncertainty estimation relies on the LLM that we use to evaluate the prompt-response pair. Therefore, it is not surprising to see that in the TriviaQA task, for answers generated by Gemma-7B, Bb-S features better uncertainty estimation than Wb-S, possibly because LLaMA2-7B, the LLM that is used as the “tool LLM” in Algorithm 1, encodes better knowledge about the uncertainty of the answers than Gemma-7B. We also note that the performance of Bb-S is not always as good as Wb-S, and we hypothesize that it is because LLMs’ output distribution differs, which could result in evaluating the uncertainty of different answers. Despite these inconsistencies, the performance of Bb-S is still strong, and these results point to a potential future avenue for estimating the uncertainty of closed-source LLMs.
@@ 4.2.2 Multiple choice (MMLU)
4.3 Interpreting the results
4.3.1 Layer comparison
For general LLMs, each token is associated with a relatively large number of hidden layers (32 layers for LLaMA2-7B for example), each of which is represented by high-dimensional vectors (4096 for LLaMA2-7B). Thus it is generally not a good practice to incorporate all hidden layers as features for the uncertainty estimation due to this dimensionality. Previous works find that the middle layer and the last layer activations of the LLM’s last token contain the most useful features for supervised learning (Burns et al., 2022; Chen et al., 2024; Ahdritz et al., 2024; Azaria and Mitchell, 2023). To investigate the layer-wise effect for uncertainty estimation, we implement our Wb-S method with features different in two aspects: (i) different layers within the LLM architecture, specifically focusing on the middle and last layers (e.g., LLaMA2-7B: 16th and 32nd layers out of 32 layers with 4096 dimensions; Gemma-7B: 14th and 28th layers out of 28 layers with 3072 dimensions); and (ii) position of token activations, including averaging hidden activations over all the prompt/answer tokens or utilizing the hidden activation of the last token. The second aspect makes sense when the output contains more than one token, so we conduct this experiment on the natural language generation tasks only. Figure 1 gives a visualization of the comparison result. While the performances of these different feature extraction ways are quite similar in terms of performance across different tasks and LLMs, activation features from the middle layer generally perform better than the last layer. This may come from the fact that the last layer focuses more on the generation of the next token instead of summarizing information of the whole sentence, as has been discussed by Azaria and Mitchell (2023).
(a) Features from Gemma-7B
(b) Features from LLaMA2-7B
4.3.2 Scaling effect
Figure 2: (Left) Using the hidden activations of LLaMA-2-7B and LLaMA-2-13B to estimate the uncertainty of the answer provided by Gemma-7B. (Right) Using the hidden activations of Gemma-2B and Gemma-7B to estimate the uncertainty of the answer provided by LLaMA-7B.
We find that there is no significant performance difference between LLaMA2-7B and LLaMA2-13B, nor between Gemma-2B and Gemma-7B (Gemma does not offer a 13B model). Our result suggests that larger LLM do not necessarily encode better knowledge about uncertainty. Furthermore, based on the performance of Wb-S in Figure 2, we suggest use the same LLM to generate the output and evaluate the uncertainty.
4.3.3 Histogram of correlations
Figure 3 plots the histograms of the pairwise correlations between the neuron activations and the labels (whether the LLM’s response is correct). We make two observations here: First, for both LLMs, some neurons have a significantly positive (or negative) correlation with the label. We can interpret these neurons as the uncertainty neuron for the corresponding task. When these neurons are activated, the LLMs are uncertain about their responses. Second, Gemma-7B has more significant neurons than LLaMA2-7B, and this is consistent with the better performance of Gemma-7B in Table 1 and Table 2. Also, this reinforces that the hidden activations of the LLMs contain uncertainty information about the LLM’s output.
Figure 3: The histograms of the pairwise correlations on the TriviaQA task between the neuron activations and the labels (whether the LLM’s response is correct), where the neural values are the last-token hidden activations of answers from the middle layer (left) and the last layer (right) of two models respectively.
Figure 4 plots some example neurons’ activation by selecting the neurons with the largest and smallest correlations in Figure 3. More such neurons can be found in Figure 6 in the Appendix. These neurons as an individual indicator exhibit different distributional patterns when the response is correct compared to when the response is incorrect, and thus reflect the uncertainty of the LLM’s responses.
(a) Neurons from the middle layer of Gemma-7B
(b) Neurons from the middle layer of LLaMA2-7B
Figure 4: Distribution of values from particular neurons on TriviaQA dataset.
4.4 Calibration performance
In Section 3.2, we distinguish the two tasks of uncertainty estimation and uncertainty calibration. Throughout the paper, we have been focused on improving the performance on the task of uncertainty estimation – to predict when the LLM is uncertain about its response. Generally, a better uncertainty estimation model leads to one with better calibration performance. The calibration (or recalibration) of the uncertainty estimation model can be indeed reduced to the classic ML setting which does not involve the LLM. Table 3 gives the calibration performance and we see an advantage of our supervised methods over benchmark methods consistent with the AUROC performance in Table 1. We adopt the histogram binning method here because we find that the temperature scaling method and the Platt scaling method will give all predicted scores concentrated within a small range such as $[0.2,0.6]$. We also do not exclude the possibility that the other calibration methods can give even better performance. The point to make here is that uncertainty estimation and uncertainty calibration are two closely related tasks. Note that (i) a better uncertainty estimation model leads to a better calibration performance and (ii) the LLMs are pre-trained and not designed for these NLP tasks in the first place (see Section 3.3) so that there is no uncertainty score readily available (as the predicted probabilities for the image classifiers); we emphasize the importance of an extra uncertainty estimation procedure as our supervised one so to extract the uncertainty information from the inside of the LLMs.
Metric Dataset Model Benchmarks Ours
Max Pro Avg Pro Max Ent Avg Ent SU A4C Bb-S Gb-S Wb-S
NLL TriviaQA G-7B 0.540 0.586 0.530 0.494 0.536 0.623 0.292 0.366 0.429
L-7B 0.486 0.459 0.508 0.569 0.530 0.618 0.482 0.455 0.414
CoQA G-7B 0.653 0.553 0.525 0.602 0.548 0.574 0.464 0.657 0.488
L-7B 0.594 0.613 0.705 0.725 0.547 0.565 0.598 0.778 0.519
WMT-14 G-7B 0.605 0.481 0.692 0.564 0.627 0.566 0.509 0.527 0.461
L-7B 0.664 0.603 0.703 0.605 0.710 0.665 0.675 0.600 0.560
ECE TriviaQA G-7B 0.031 0.050 0.039 0.045 0.031 0.036 0.033 0.037 0.044
L-7B 0.055 0.026 0.044 0.045 0.060 0.021 0.042 0.059 0.052
CoQA G-7B 0.056 0.046 0.059 0.068 0.056 0.027 0.049 0.065 0.032
L-7B 0.038 0.036 0.045 0.046 0.035 0.025 0.045 0.049 0.022
WMT-14 G-7B 0.046 0.042 0.075 0.030 0.043 0.042 0.040 0.042 0.037
L-7B 0.037 0.026 0.033 0.021 0.033 0.027 0.051 0.041 0.039
Brier TriviaQA G-7B 0.162 0.168 0.163 0.149 0.177 0.215 0.071 0.120 0.076
L-7B 0.148 0.150 0.152 0.161 0.160 0.207 0.144 0.145 0.131
CoQA G-7B 0.153 0.155 0.161 0.161 0.160 0.167 0.147 0.158 0.143
L-7B 0.174 0.178 0.178 0.183 0.182 0.189 0.173 0.181 0.161
WMT-14 G-7B 0.184 0.157 0.213 0.156 0.204 0.195 0.177 0.159 0.149
L-7B 0.229 0.210 0.234 0.212 0.236 0.236 0.227 0.208 0.192
@@ 4.5 Transferability
In this subsection, we evaluate the robustness of our methods under the out-of-distribution (OOD) setting for both question-answering and multiple-choice tasks.
@@
G-7B TriviaQA 0.956(0.964) 0.837(0.893) 0.899(0.963) 0.797 0.712
CoQA 0.707(0.747) 0.706(0.729) 0.722(0.769) 0.720 0.657
L-7B TriviaQA 0.777(0.856) 0.848(0.852) 0.860(0.879) 0.835 0.817
CoQA 0.708(0.706) 0.702(0.713) 0.717(0.751) 0.699 0.647
Table 4: Transferability of the trained uncertainty estimation model across different groups of subjects in MMLU and question-answering datasets. For our proposed Bb-S, Gb-S, and Wb-S methods, values within the parentheses $(\cdot)$ represent the AUROCs where the calibration model is trained and tested on the same group of subjects or dataset, while values outside the parentheses are AUROCs where the model is trained on another group of subjects or dataset. The Best GB and Best BB columns refer to the best AUROC achieved by the unsupervised grey-box baselines and black-box baselines (fully listed in Table 1 and Table 2), respectively.
Table 4 summarizes the performance of these OOD experiments. As expected, for all the methods, there is a slight drop in terms of performance compared to the in-distribution setting (reported by the numbers in the parentheses in the table). We make the following observations based on the experiment results. First, based on the performance gap between in-distribution and OOD evaluation, it is evident that although incorporating white-box features such as hidden activations makes the model more susceptible to performance decreases on OOD tasks, these features also enhance the uncertainty estimation model’s overall capacity, and the benefits outweigh the drawbacks. It is also noteworthy that even in these scenarios of OOD, our Wb-S and Bb-S method almost consistently outperform baseline approaches. Overall, the robustness of our methods shows that the hidden layers’ activations within the LLM exhibit similar patterns in encoding uncertainty information to some extent. The performance drop (from in-distribution to OOD) observed in the MMLU dataset is notably less than that in the question-answering dataset, which may stem from the larger disparity between the CoQA and TriviaQA datasets compared to that between two distinct groups of subjects within the same MMLU dataset. This suggests that in cases of significant distributional shifts, re-training or re-calibrating the uncertainty estimation model using test data may be helpful.

### v2

@@ 3.4 Three regimes of supervised uncertainty estimation
A natural implementation of our supervised learning approach involves using an LLM to generate the response $\bm{y}$ for input $\bm{x}$, and extracting insights on confidence from the same LLM’s hidden layers’ activations. This method functions effectively with white-box LLMs where hidden activations are accessible. Note that obtaining hidden layers’ activations merely requires an LLM and the prompt-response pair $(\bm{x},\bm{y})$. Therefore, it is not mandatory for $(\bm{x},\bm{y})$ to be generated by the LLM that provides the hidden layers’ activations. Based on this observation, we note that the extra knowledge of uncertainty can come from the hidden layers of any white-box LLM that takes as input the $(\bm{x},\bm{y})$ pair, not necessarily from the LLM that generates $(\bm{x},\bm{y})$. It is natural to start with the belief that the LLM which generates $(\bm{x},\bm{y})$ should have more information about the uncertainty of $\bm{y}$. However, any white-box LLM can output the hidden activations corresponding to $(\bm{x},\bm{y})$.
@@
Grey-box supervised uncertainty estimation (Gb-S): This Gb-S regime constructs the features only from the grey-box source, that is, those features relying on the probability and the entropy (such as those in Table 5 in Appendix C) but it ignores the hidden-layer activations.
@@ 4.1 LLMs, tasks, benchmarks, and performance metrics
LLMs. For our numerical experiments, we mainly consider three open-source LLMs, LLaMA2-7B (Touvron et al., 2023), LLaMA3-8B (AI@Meta, 2024) and Gemma-7B (Gemma Team et al., 2024) as $p_{\theta}$ defined in Section 2. For certain experiments, we also employ the models of LLaMA2-13B and Gemma-2B. We also use their respective tokenizers as provided by Hugging Face. We do not change the parameters/weights $\theta$ of these LLMs.
Tasks and Datasets. We mainly consider three tasks for uncertainty estimation, question answering, multiple choice, and machine translation. All the labeled datasets for these tasks are in the form of $\{(\bm{x}_{i},\bm{y}_{i,\text{true}})\}_{i=1}^{n}$ where $\bm{x}_{i}$ can be viewed as the prompt for the $i$-th sample and $\bm{y}_{i,\text{true}}$ the true response. We adopt the few-shot prompting when generating the LLM’s response $\bm{y}_{i}$, and we use 5 examples in the prompt of the multiple-choice task and 3 examples for the remaining natural language generation tasks. This enables the LLM’s in-context learning ability (Radford et al., 2019; Zhang et al., 2023) and ensures the LLM’s responses are in a desirable format. We defer more details of the few-shot prompting to Appendix C.1. The three tasks are:
@@
[…]
@@
[…]
Our methods. We follow the discussions in Section 3.4 and implement three versions of our proposed supervised approach: black-box supervised (Bb-S), grey-box supervised (Gb-S), and white-box supervised (Wb-S). These models have the same pipeline of training the uncertainty estimation model and the difference is only on the availability of the LLM. For the Bb-S method, we use the Gemma-7B as the model $q_{\theta}$ to evaluate the uncertainty of LLaMA2-7B/LLaMA3-8B $p_{\theta}$ (treated as a black-box), and reversely, use LLaMA2-7B to evaluate Gemma-7B. The supervised uncertainty model $\hat{g}$ is trained based on the random forest model (Breiman, 2001). Details on the feature construction and the training of the random forest model are deferred to Appendix C.2.
Performance metrics. For the model evaluation, we follow Filos et al. (2019); Kuhn et al. (2023) and compare the performance of our methods against the benchmark using the generated uncertainty score to predict whether the answer is correct. The area under the receiver operator characteristic curve (AUROC) metric is employed to measure the performance of the uncertainty estimation. As discussed in Section 3.2, AUROC works as a good metric for the uncertainty estimation task whereas for the uncertainty calibration task, we follow the more standard calibration metrics and present the results in Section B.
@@ 4.2.1 Question answering and machine translation
MaxL AvgL MaxE AvgE SU A4C Bb-S Gb-S Wb-S
TriviaQA G-7B 0.857 0.862 0.849 0.854 0.847 0.534 0.879 0.866 0.882
L-7B 0.565 0.761 0.761 0.773 0.678 0.526 0.925 0.811 0.897
L-8B 0.838 0.851 0.849 0.853 0.826 0.571 0.843 0.861 0.874
CoQA G-7B 0.710 0.708 0.725 0.708 0.674 0.515 0.737 0.737 0.762
L-7B 0.535 0.600 0.603 0.580 0.541 0.502 0.848 0.667 0.807
L-8B 0.692 0.697 0.716 0.699 0.684 0.506 0.745 0.737 0.769
WMT-14 G-7B 0.668 0.589 0.637 0.811 0.572 0.596 0.863 0.829 0.855
L-7B 0.606 0.712 0.583 0.711 0.513 0.506 0.792 0.724 0.779
L-8B 0.554 0.685 0.616 0.729 0.510 0.502 0.700 0.724 0.745
@@
Table 1: Out-of-sample AUROC performance for benchmarks and our methods on natural language generation tasks. G-7B, L-7B, and L-8B represent Gemma-7B, LLaMA2-7B, and LLaMA-8B, respectively. The columns MaxL, AvgL, MaxE, and AvgE all come from Manakul et al. (2023). The column SU implements the semantic uncertainty estimation by Kuhn et al. (2023), and the column A4C implements the ask-for-confidence method by Tian et al. (2023). The columns Bb-S, Gb-S, and Wb-S represent respectively the three regimes (black-box supervised, grey-box supervised, and white-box supervised) of our supervised method with details in Section 3.4.
@@
We make several remarks on the numerical results. First, our methods generally have a better performance than the existing benchmarks. Note that the existing benchmarks are mainly unsupervised and based on one single score, and also that our method proceeds with the most standard pipeline for supervised training of an uncertainty estimation model. The advantage of our method should be attributed to the supervised nature and the labeled dataset. While these unsupervised benchmark methods can work in a larger scope than these NLP tasks (though they have not been extensively tested on open questions yet), our methods rely on the labeled dataset. But in addition to these better numbers, the experiment results show the potential of labeled datasets for understanding the uncertainty in LLM’s responses. In particular, our method Gb-S uses features including the benchmark methods, and it shows that some minor supervised training can improve a lot upon the ad-hoc uncertainty estimation based on one single score such as MaxL or MaxE.
@@
Second, our method Wb-S has a clear advantage over our method Gb-S. Note that these two methods differ in that the Wb-S uses the hidden activations while the Gb-S only uses probability-related (and entropy-related) features. This implies that the hidden activations do contain uncertainty information which we will investigate more in Appendix A. Also, we note from the table that there is no single unsupervised grey-box method (under the Benchmarks columns) that consistently surpasses others across different datasets/NLP tasks. For example, among all these unsupervised benchmark methods for grey-box LLMs, AvgE emerges as a top-performing one for the Gemma-7B model when applied to the machine translation task, but it shows the poorest performance for the same Gemma-7B model when tested on the question-answering CoQA dataset. This inconsistency highlights some caveats when using the unsupervised approach for uncertainty estimation of LLMs.
@@
Lastly, we note that the Bb-S method has a similar or even better performance as the Wb-S method. As discussed in Section 3.4, the performance of uncertainty estimation relies on the LLM that we use to evaluate the prompt-response pair. Therefore, it is not surprising to see that in the question-answering task, for answers generated by LLaMA2-7B, Bb-S features better uncertainty estimation than Wb-S, possibly because Gemma-7B, the LLM that is used as the “tool LLM” in Algorithm 1, encodes better knowledge about the uncertainty of the answers than LLaMA-7B. We also note that the performance of Bb-S is not always as good as Wb-S, and we hypothesize that it is because LLMs’ output distribution differs, which could result in evaluating the uncertainty of different answers. Despite these inconsistencies, the performance of Bb-S is still strong, and these results point to a potential future avenue for estimating the uncertainty of closed-source LLMs.
@@ 4.2.2 Multiple choice (MMLU)
We defer more numerical experiments and visualization to Appendices A and B where we investigate more on (i) the effect of the choice of layers; (ii) the scale of the LLMs used; (iii) the uncertainty neurons of the LLMs; and (iv) the calibration performance.
@@ 4.3 Transferability
In this subsection, we evaluate the robustness of our methods under the OOD setting.
@@
G-7B TriviaQA 0.842(0.879) 0.861(0.866) 0.861(0.882) 0.862 0.847
CoQA 0.702(0.737) 0.722(0.737) 0.730(0.762) 0.725 0.674
L-7B TriviaQA 0.917(0.925) 0.801(0.811) 0.881(0.897) 0.773 0.678
CoQA 0.825(0.848) 0.623(0.667) 0.764(0.807) 0.603 0.541
L-8B TriviaQA 0.813(0.843) 0.859(0.861) 0.863(0.874) 0.853 0.826
CoQA 0.710(0.745) 0.714(0.737) 0.725(0.769) 0.716 0.684
Table 3: Transferability of the trained uncertainty estimation model across different groups of subjects in MMLU and question-answering datasets. For our proposed Bb-S, Gb-S, and Wb-S methods, values within the parentheses ()($\cdot$)( ⋅ ) represent the AUROCs where the uncertainty estimation model is trained and tested on the same group of subjects or dataset, while values outside the parentheses represent models trained on another group of subjects or dataset. The Best GB and Best BB columns refer to the best AUROC achieved by the unsupervised grey-box baselines and black-box baselines (fully listed in Table 1 and Table 2), respectively.
Table 3 summarizes the performance of these OOD experiments. As expected, for all the methods, there is a slight drop in terms of performance compared to the in-distribution setting (reported by the numbers in the parentheses in the table). We make the following observations based on the experiment results. First, based on the performance gap between in-distribution and OOD evaluation, it is evident that although incorporating white-box features such as hidden activations makes the model more susceptible to performance decreases on OOD tasks, these features also enhance the uncertainty estimation model’s overall capacity, and the benefits outweigh the drawbacks. It is also noteworthy that even in these scenarios of OOD, our Wb-S and Bb-S method almost consistently outperform corresponding baseline approaches. Overall, the robustness of our methods shows that the hidden layers’ activations within the LLM exhibit similar patterns in encoding uncertainty information to some extent. The performance drop (from in-distribution to OOD) observed in the MMLU dataset is notably less than that in the question-answering dataset, which may stem from the larger disparity between the CoQA and TriviaQA datasets compared to that between two distinct groups of subjects within the same MMLU dataset. This suggests that in cases of significant distributional shifts, re-training or re-calibrating the uncertainty estimation model using test data may be helpful.